x + 2y = 3
x = 5 - 2y
The solution is what ordered pair (?,?)

Answer by jessica43(140)

I think you may have written this problem down incorrectly:
x + 2y = 3
x = 5 - 2y
Plug the second into the first:
x + 2y = 3
(5-2y) + 2y = 3
5=3
This is not possible.....